# Enhancing multimodal analogical reasoning with Logic Augmented Generation

### v1

@@ VI-D4 Metaphor detection and understanding: comparison with open-source LLM
We compare the best performing results of metaphor detection with Llama, 3.1 70B Instruct.
@@ VII-A Application of LAG to multimodal metaphor detection and interpretation
The results outlined in Section V show that for the task of metaphor detection, the augmented KG according to a prompt and a base graph that constrain the LLM to adhere to the Blending Ontology outperforms the current metaphor detection and understanding baselines. This approach can be used to extract knowledge graphs from unstructured multimodal metaphorical data. As such, it could have effects in different applications, for example, improving hate speech, which often has metaphorical features [60], and analyzing debates in the press and online media. For what concerns visual metaphors, since humans tend to struggle to identify them, the automated understanding of visual blends, for example in advertisement, can bridge this gap. Additionally, using an open-source LLM like Llama can decrease the performance concerning detection, but not understanding tasks. Ablation studies show that while removing the graph reduces the performance in all cases, removing the blending heuristics in the prompting reduces it in all cases besides visual metaphor understanding. While this is a slight difference, it could be motivated by the fact visual metaphors can be wrongly described in text before processing. Another explanation for this slight performance drop is shown by the error analysis (see Section VII-B), where incorrect property was the most frequent issue both for humans [32] and LLMs: the objects are detected in the images, but there is still a struggle with extracting the correct properties. Because half of the images are advertisements, which typically present clearer context, the intended properties of the objects are usually clearly defined. By not specifying the purpose of the image and its context, we make it challenging for both humans and machines to accurately determine the metaphorical property via a blending mechanism that tends to be more open-ended without contextual specifications.
@@ VII-B Quality of the augmented graphs and error analysis
We provide a sample of 16 generated graphs, 8 derived from correctly identified metaphorical sentences, 8 derived from visual metaphors, to two knowledge engineers experts in conceptual blending and conceptual metaphor theories. We evaluate the representation’s adequacy, correctness, and completeness on a Likert scale from 1 to 7 concerning detected property, source, target and element mappings. The results highlight a consistent pattern where textual components generally outperform their visual counterparts across all evaluated metrics. Specifically, correctness and adequacy in text scored at 5.1 and 6.3, suggesting a positive assessment, whereas the visual aspects scored slightly lower at 4.9 and 5 respectively, indicating room for improvement. The completeness of the text was rated at 5.2, while the visual completeness was lower at 4.4, showing a need for better integration or representation of visual elements to match the textual analysis. Overall, these results suggest a positive performance with a disparity between text and visual outputs, probably because the chosen examples were the ones taken by Petridis et al. [32] and they weren’t further annotated with an RDF syntax to be consistent with the human evaluation.
VII-B1 Error analysis
While metaphor detection shows good performance in known datasets, for conceptual metaphor understanding, errors of the best performing LAG configuration follow the categorization proposed in Wachowiak and Gromann  [40] with the addition of a new category found, the incorrect switching of source and target with respect to the gold standard annotation. Furthermore, generic conceptual metaphors tend to be detected more often than scientific ones.
@@ VII-B1 Error analysis
In particular, misalignment in conceptual metaphor mapping is predominantly due to wrong subelement mapping, which accounts for approximately 56.5% in the WG dataset and 57.1% in the BCTMD. Annotations that are too specific contribute significantly as well, with 23.6% and 28.6% of errors in WG and BCTMD respectively, while overly general descriptions, appearing in 9.3% of WG and 14.3% of BCTMD cases, showing difficulty in identifying the correct taxonomical level of the metaphor. Additionally, less frequent errors, such as switching the source and target (3.24% in WG) and identifying as literal something that was a metaphor (6.94% in WG), show further layers of complexity in the metaphor mapping process, such as misunderstanding the role of source and target domains and determining the metaphoricity level (a sentence might seem literal when the perceived conventionality of the metaphor is high).
@@
These results underscore the need to train and test automatic methods on domain-specific figurative datasets, but also that analogical reasoning is generally not properly performed in these tasks. On the other hand, visual metaphor understanding shows different kinds of errors. The error analysis still follows the one done with human participants in Petridis et al.  [32], with errors involving incorrect objects (21% of the errors). For example, a LLM will identify the shape of a gun as a whistle, or pink Nike shoes as fishes, usually skewing the interpretation. The most common error, however, is incorrect property (57% of the errors): if the LLM correctly identifies the objects, source and target, but still misinterprets the property. For example, in the case of the blending of the car keys and the gun (displayed in Figure LABEL:fig:visualmet), with the annotated property being dangerous, the predicted property is powerful. This error is also observed in human participants and tied to their values and beliefs over the objects displayed in the blend, a further area of investigation. Finally, in 21% of cases, even when the LLM identifies source, target and connecting property, the role mapping explanation can show a misunderstanding of the metaphor (incorrect target symbol). Thus, the evaluation of the property alone as proposed in previous studies [32] is not sufficient to see if the system has understood a metaphor.
@@
VII-C Application of LAG to metaphor generation
In addition to detecting existing metaphors, our framework offers a structured approach to generating new effective metaphors. By formalizing the process of blending conceptual frames, the system can suggest novel metaphors by creatively reconfiguring known concepts into new, plausible configurations according to a specified ontology. This aspect of the framework holds particular promise for applications in creative AI fields, such as advertising, literature, and automated content generation. In future research, we aim to develop a method for generation that addresses limitations found in other studies, such as that described in Chakrabarty et al.  [61]. The authors in this study prompt models to generate visual metaphors using generic, implicit elements, but do not specify how these should be effectively combined to produce meaningful metaphors. Indeed, one reported error in the study was the omission of the target domain or essential elements that were not explicitly requested in the image. Finding the LLM’s errors in understanding visual metaphors can thus improve its ability to generate metaphors in a closer way to our thought processes. This enrichment can foster narrative creation, offering fresh and impactful ways to communicate complex ideas or emotions, and enhancing the engagement and relatability of AI-generated content. Furthermore, with the human-validated data about LLM-generated overarching concepts, scores information can provide some input for the finetuning of weighted blending representations.
VII-D Interpretability of metaphorical datasets
@@
VII-D1 Context and culture
Our research demonstrates that annotating the source and target elements in computational metaphor analysis is insufficient for fully grasping metaphorical meaning. For instance, as illustrated in Figure LABEL:fig:visualmet, the LLM accurately identifies all objects within the image that form the metaphor. However, although the blending property is annotated as “dangerous”, the LLM interprets it as “power”. In the absence of external context that can further frame the reference elements needed for comprehension, both interpretations are arguably valid.
This and other examples from our experiments reveal two key insights. First of all, the perspective adopted by an LLM in interpreting a metaphor reflects its underlying cultural background. In this case, the dual perception of a gun as either powerful or dangerous underscores how cultural context shapes interpretation.
Furthermore, for systems (and users) to reliably converge on metaphorical annotated properties, providing additional context is essential. This requirement for context may also help explain the relatively poor performance observed in human responses, such as distinguishing whether an advertisement is a public campaign, a comic strip, or a standard advertisement.
VII-D2 The dynamics of metaphor annotation
Beyond culture and context-dependence, the human annotation showed that some sentences’ gold standard metaphors in the WG, BCTMD and the Visual Metaphors dataset had multiple possible correct answers or even multiple answers in the same sentence. For example, for the sentence “This mini-controversy erupted when Republicans introduced a string of amendments in a final effort to obstruct passage of the reconciliation bill” the gold-standard metaphor is inhibitting enactment of legislation is impeding motion, while the predicted one is controversy is volcano eruption, which has been noted as correct by the evaluators but would be deemed as incorrect with automatic assessment.
This phenomenon shows the need to keep both the human and the metaphor interpretation process flexible. However, current annotated metaphorical datasets only contain one possible gold standard answer. Future work should address the possibility to incorporate multiple correct answers for a reference text.
@@ VII-E Limitations
Despite its strengths, the current framework has possible improvements. Experiments are limited to English; more text and image data are needed for a more comprehensive assessment. While the work on conceptual metaphors shows some correlation with automated measures, manual evaluation remains essential. Indeed, another survey on the quality of the augmented graphs with a higher quantity of graphs and participants is needed. Additionally, the model’s computational complexity and resource demands may cause time constraints with large datasets or real-time applications. Ongoing refinement is thus necessary to improve scalability and efficiency.
VIII Conclusion and Future Work
In this paper, we presented a Logic-Augmented Generation-based framework that automatically generates extended knowledge graph triples from a base graph using ontology prompting heuristics to enable effective analogical reasoning on various types of unlabeled figurative data. We gathered four datasets for metaphor detection and understanding and introduced a novel dataset featuring domain-specific scientific conceptual metaphors. Our results in multimodal metaphor detection and understanding outperform current baselines with Claude Sonnet 3.5 and even surpass human performance in comprehending visual metaphors, although the accuracy for scientific metaphors remains significantly lower than for generic ones. The error analysis confirms recent claims that, while LLMs excel at surface associations, they still lack robust relational reasoning processes, partly due to insufficient contextual explanations in existing datasets. Despite this shortcoming, which calls for more diverse and fine-grained metaphorical data, leveraging the Blending Ontology to generate augmented knowledge graphs enhances output explainability compared to current methods, which often fail to clarify why, even when the objects of a metaphor are detected, the metaphorical property is not correctly identified. The proposed approach therefore advances computational models for multimodal figurative language understanding in unstructured data, with significant implications for AI in natural language processing, creativity, and applications such as hate speech detection. Future work will focus on refining the framework’s computational complexity and testing it on a broader range of annotated data.

### v2

@@ VI-D4 Metaphor detection and understanding: comparison with open-source LLM
In Table VI, we compare the Claude-based best performing results of metaphor detection with the ones from Llama, 3.1 70B Instruct, showing Claude outperforms the open-source alternative.
@@ VI-D5 Ablation studies
VI-D6 Quality of the augmented graphs and error analysis
The results of the survey provided to two knowledge engineers on the quality of the augmented graphs highlight a consistent pattern where textual components generally outperform their visual counterparts across all evaluated metrics. Specifically, on a Likert scale from 1 to 7, correctness and adequacy in text scored at 5.1 and 6.3, suggesting a positive assessment, whereas the visual aspects scored slightly lower at 4.9 and 5 respectively, indicating room for improvement. The completeness of the text was rated at 5.2, while the visual completeness was lower at 4.4, showing a need for better integration or representation of visual elements to match the textual analysis.
For what concerns the error analysis, misalignment in conceptual metaphor mapping is predominantly due to wrong subelement mapping, which accounts for approximately 56.5% in the WG dataset and 57.1% in the BCTMD. Annotations that are too specific contribute significantly as well, with 23.6% and 28.6% of errors in WG and BCTMD respectively, while overly general descriptions appear in 9.3% of WG and 14.3% of BCTMD cases. Less frequent errors concern switching the source and target (3.24% in WG) and identifying as literal something that was a metaphor (6.94% in WG). For visual metaphor understanding, the error analysis still follows the one done with human participants in Petridis et al. [52], with errors mostly involving incorrect objects (21% of the errors) and incorrect property (57% of the errors). In 21% of cases, even when the LLM identifies source, target and connecting property, the role mapping explanation can show a misunderstanding of the metaphor (incorrect target symbol). For a more detailed analysis along with example errors, see the Supplementary Material.
@@ VII-A Application of LAG to multimodal metaphor detection and interpretation
The results outlined in Section V show that for the task of metaphor detection, the proposed LAG-based framework outperforms the current metaphor detection and understanding baselines. This approach can be used to extract knowledge graphs from unstructured multimodal metaphorical data. As such, it could have effects in different applications, for example, improving hate speech, which often has metaphorical features [70], and analyzing debates in the press and online media. For what concerns visual metaphors, since humans tend to struggle to identify them [52], the automated understanding of visual blends can bridge this gap. Additionally, using an open-source LLM like Llama can decrease the performance concerning detection, but not understanding tasks.
Ablation studies show that while removing the graph reduces the performance in all cases, removing the blending heuristics in the prompt reduces it in all cases besides visual metaphor understanding. While this is a slight difference, it could be motivated by the fact that visual metaphors can be wrongly described in text before processing. Another explanation for this slight performance drop is shown by the error analysis (see Section VII-B), where incorrect property was the most frequent issue both for humans [52] and LLMs: the objects are detected in the images, but there is still a struggle with extracting the correct properties. Because half of the images are advertisements, which typically present clearer context, the intended properties of the objects are usually clearly defined. By not specifying the purpose of the image and its context, we make it challenging for both humans and machines to accurately determine the metaphorical property via a blending mechanism that tends to be more open-ended without contextual specifications. Additionally, the low accuracy score observed in detecting scientific metaphors appears to reflect inherent limitations in current LLMs rather than a specific weakness of the LAG approach. Scientific metaphors often require domain-specific context and specialized knowledge that may not be well-represented in LLM training data. Also, scientific metaphors show multiple levels of conventionality, tending to result more literal as they come to be widespread in language.
@@ VII-B Quality of the augmented graphs and error analysis
The results from the questionnaire to analyze the quality of the augmented graphs suggest a positive performance with a disparity between text and visual outputs, probably because the chosen examples were the ones taken by Petridis et al. [52] and they weren’t further annotated with an RDF syntax in order to be consistent with the human evaluation.
@@
For what concerns error analysis, overly general descriptions show difficulty in identifying the correct taxonomical level of the metaphor. Additionally, less frequent errors, such as switching the source and target and identifying as literal something that was a metaphor, show further layers of complexity in the metaphor mapping process, such as misunderstanding the role of source and target domains and determining the metaphoricity level (a sentence might seem literal when the perceived conventionality of the metaphor is high).
@@
These results underscore the need to train and test automatic methods on domain-specific figurative datasets, but also that analogical reasoning is generally not properly performed in these tasks. On the other hand, visual metaphor understanding shows different kinds of errors. Errors involving incorrect objects indicate poor image understanding skills, for example, a LLM will identify the shape of a gun as a whistle, or pink Nike shoes as fishes, usually skewing the interpretation. The most common error, however, is incorrect property: if the LLM correctly identifies the objects, source and target, but still misinterprets the property. For example, in the case of the blending of the car keys and the gun (displayed in Figure 4), with the annotated property being dangerous, the predicted property is powerful. This error is also observed in human participants and tied to their values and beliefs over the objects displayed in the blend, a further area of investigation. Furthermore, the presence of the incorrect target symbol error shows that the evaluation of the property alone as proposed in previous studies [52] is not sufficient to verify if the system has understood a metaphor.
@@
VII-C Interpretability of metaphorical datasets
@@
VII-C1 Context and culture
Our research demonstrates that annotating the source and target elements in computational metaphor analysis is insufficient for fully grasping metaphorical meaning. For instance, as illustrated in Figure 4, the LLM accurately identifies all objects within the image that form the metaphor. However, although the blending property is annotated as dangerous, the LLM interprets it as powerful. In the absence of external context that can further frame the reference elements needed for comprehension, both interpretations are arguably valid.
This and other examples from our experiments reveal two key insights. First of all, the perspective adopted by a LLM in interpreting a metaphor reflects its underlying cultural background. Furthermore, for systems (and users) to reliably converge on metaphorical annotated properties, providing additional context is essential. This requirement for context may also help explain the relatively poor performance observed in human responses, such as distinguishing whether an advertisement is a public campaign, a comic strip, or a standard advertisement.
VII-C2 The dynamics of metaphor annotation
Beyond culture and context-dependence, the human annotation showed that some sentences’ gold standard metaphors in the WG, BCTMD and the Visual Metaphors dataset had multiple possible correct answers or even multiple answers in the same sentence. For example, for the sentence “This mini-controversy erupted when Republicans introduced a string of amendments in a final effort to obstruct passage of the reconciliation bill” the gold-standard metaphor is inhibiting enactment of legislation is impeding motion, while the predicted one is controversy is volcano eruption, which has been noted as correct by the evaluators but would be deemed as incorrect with automatic assessment. This phenomenon shows the need to keep both the human and the metaphor interpretation process flexible. However, current annotated metaphorical datasets only contain one possible gold standard answer. Future work should address the possibility to incorporate multiple correct answers for a reference text.
VII-D Future application of LAG to metaphor generation
In addition to detecting existing metaphors, our framework offers a structured approach to generating new effective metaphors. By formalizing the process of blending conceptual frames, the system can suggest novel metaphors by creatively reconfiguring known concepts into new, plausible configurations according to a specified ontology. This aspect of the framework holds particular promise for applications in creative AI fields, such as advertising, literature, and automated content generation. In future research, we aim to develop a method for generation that addresses limitations found in other studies, such as the ones described in previous studies [71, 42]. Specifically, the authors in Lieto et al. [42] note that top-down frame taxonomie must be complemented by more flexible generators.Chakrabarty et al. prompt models to generate visual metaphors through generic, implicit elements, but do not specify how these should be effectively combined to produce meaningful metaphors. Indeed, one reported error in the study was the omission of the target domain or essential elements that were not explicitly requested in the output image. Finding the LLM’s pitfalls in understanding visual metaphors can thus improve its ability to generate metaphors in a closer way to our thought processes. This enrichment can foster narrative creation, offering fresh and impactful ways to communicate complex ideas or emotions, and enhancing the engagement and relatability of AI-generated content.
@@ VII-E Limitations
Despite its strengths, the current framework has possible improvements. Experiments are limited to English; more multimodal data is needed for a more comprehensive assessment. While the work on conceptual metaphors shows some correlation with automated measures, manual evaluation remains essential. Indeed, another survey on the quality of the augmented graphs with a higher quantity of graphs and expert participants is needed. Additionally, the model’s computational complexity and resource demands may cause time constraints with large datasets or real-time applications. Ongoing refinement is thus necessary to improve scalability and efficiency. For what concerns data leakage, while legacy textual benchmarks (e.g., MOH-X, TroFi) may already reside in the analyzed LLMs’ web-scale pre-training corpus, the BTCMD corpus and the visual-metaphor dataset were newly curated and remained unpublished online before evaluation, ensuring no train–test leakage for these two resources.
VIII Conclusion
In this paper, we introduced a Logic-Augmented Generation pipeline that turns multimodal data into extended knowledge graphs by injecting a Blending Ontology at prompt time. Tested on four metaphor benchmarks, the system outperformed baselines by up to +12 $F_{1}$ on text and +8 accuracy on visual metaphors, and even surpassed the human gold standard on the visual metaphors dataset. Performance on scientific metaphors, however, trailed generic tasks by 17 percentage points, showing that domain-specific metaphors need to be treated differently from generic ones.
The error analysis supports recent findings that today’s LLMs excel at surface associations but falter on relational reasoning; many failures stem from sparse or mismatched contextual cues in existing corpora. Despite this shortcoming, which calls for more diverse and fine-grained metaphorical data, leveraging the Blending Ontology to generate augmented knowledge graphs enhances output explainability compared to current methods, which often fail to clarify why, even when the objects of a metaphor are detected, the metaphorical property connecting the two is not correctly identified. The proposed approach advances computational models for multimodal figurative language understanding in unstructured data, with significant implications for AI in natural language processing, creativity, and applications such as hate speech detection.